I'm stuck on a problem could someone please help me?
Find the value for"a" if the points are the indicated apart.
(2,a) and (-3,-2)when d=radical sign 34

Answer by stanbon(48568)

Hopefully you know the distance formula
d^2=[(difference of x)^2 +(diff of y's)^2]
(2+3)^2+(a+2)^2 = 34
25 +(a+2)^2=34
(a+2)^2=9
a+2=3 or a+2=-3
a=1 or a=-5
Cheers,
Stan H.